# Ten Years of ZMap

### v1

@@ Abstract
Since ZMap’s debut in 2013, networking and security researchers have used the open-source scanner to write hundreds of research papers that study Internet behavior. In addition, ZMap powers much of the attack-surface management and security ratings industries, and more than a dozen security companies have built products on top of ZMap. Behind the scenes, much of ZMap’s behavior—ranging from its pseudorandom IP generation to its packet construction—has quietly evolved as we have learned more about how to scan the Internet. In this work, we quantify ZMap’s adoption over the ten years since its release, describe its modern behavior (and the measurements that motivated those changes), and offer lessons from releasing and maintaining ZMap.
@@ 1. Introduction
In 2013, Durumeric et al. released ZMap (durumeric2013zmap, ), an open-source network scanner that made it dramatically easier to scan the entire IPv4 address space. Since then, more than 300 research papers have used ZMap to uncover protocol flaws (adrian2015imperfect, ; aviram2016drown, ), shed light on the WebPKI, reverse engineer mercenary spyware (marczak2014governments, ), understand headline events like Heartbleed (durumeric2014matter, ) and Mirai (antonakakis2017understanding, ), and more. Beyond research, more than a dozen security companies have developed products on top of ZMap to continuously monitor companies’ attack surfaces and their third-party dependencies. At the same time, like many security tools, ZMap has also been adopted by attackers to identify vulnerable systems. In aggregate, ZMap accounts for over one-third of all Internet scan traffic today.
Since its initial release, ZMap has greatly evolved as we have learned more about how to scan the Internet. Yet, with the exception of Adrian et al.’s write up of ZMap’s 10GbE re-architecture (adrian2014zippier, ), the project team has not documented these changes in the research literature. Indeed, of the improvements Adrian et al. introduced, both the lock-free randomization algorithm and fast packet transmission approach have changed. We have also learned a great deal about how to (and not to) build measurement tools.
Motivated by requests from the community to document what we learned (claffy2019workshop, ), in this work, we present a retrospective analysis of ZMap. We cover how ZMap has been adopted (Section 2), advances in Internet scanning that challenged the tool’s initial design assumptions (Section 3), significant changes to ZMap’s behavior (Section 4), and what we learned maintaining and improving ZMap (Section 5). We hope that by analyzing this endeavor, we can help the researchers who create the next generation of Internet measurement tools.
@@ 2. Usage and Internet Landscape
Since ZMap’s release in 2013, its adoption has steadily increased (durumeric2014internet, ; anand2023aggressive, ). Today, over 33% of all IPv4 scan traffic can be fingerprinted as coming from ZMap. In this section, we describe ZMap’s adoption by academic researchers, security companies, and malicious actors.
Figure 1. ZMap-Attributed Scan Traffic—ZMap growth has accelerated significantly since 2020. In Q1 2024, 35% of global IPv4 TCP scan traffic (by packet) came from ZMap.
@@ 2.1. Empirical Analysis
One year after ZMap’s release, in 2014, Durumeric et al. found that there had been relatively little adoption and the tool was primarily being used to study academically interesting protocols like HTTP(S) (durumeric2014internet, ). More recently, in 2021, Anand et al. noted that many “aggressive” scans were from ZMap or Masscan (anand2023aggressive, ). Using the same methodology as these two studies (durumeric2014internet, ; anand2023aggressive, ), we measure ZMap adoption by analyzing scans that target at least 10 IPs in the ORION Network Telescope (orion, ). Our analysis is limited to TCP scans, since ORION does not tag scanning tools for ICMP and UDP flows.
ZMap usage has increased dramatically over the past four years (Figure 1). From January 1 to March 31, 2024, 35.4% of all IPv4 TCP scan packets originated from ZMap. While it is possible that ICMP and UDP scans use other tools, the vast majority of all scan traffic is TCP-based and ZMap still accounts for 33% of all scan packets with only TCP traffic attributed. ZMap scans follow a different distribution of ports than global scans (Figures 2 and 3). For example, ZMap accounts for only 12% of TCP/23 (Telnet) but 69% of TCP/80 and 73% of TCP/8080. In the most extreme case, 99.5% of traffic to TCP/8728 (MikroTik router API) is from ZMap, driving it to the sixth most scanned port globally. There are also dramatic regional differences in ZMap adoption (Figure 4). For example, while ZMap accounts for more than 66% of scan packets from U.S. hosts, less than 0.5% of Russian scan packets are from ZMap. The outsized use in the U.S. is largely due to adoption by American security companies (Section 2.3).
Figure 2. All TCP Scans (Top 30 Ports by Packet)
Figure 3. Top Ports Scanned by ZMap (by Packet)
@@ 2.2. Academic Research
ZMap has been used for a vast range of research purposes, from compromising RSA keys through transient faults  (sullivan2022open, ) to measuring NAT64 deployment (hsu2024first, ). To understand the studies that ZMap has enabled, we manually analyzed 1,034 papers that cite or reference ZMap through April 2024 via Google Scholar. We exclude dissertations from our analysis (since these are often comprised of published papers) as well as studies that use Censys (durumeric2015search, ). In total, we identified 307 research papers directly based on ZMap data.
While ZMap is a general measurement tool, it has most prominently been used by the security community. Notably, ZMap has been used in 38 papers to uncover protocol weaknesses in TLS and underlying cryptographic primitives (beurdouche2015messy, ; adrian2015imperfect, ; aviram2016drown, ) as well as to uncover deployment challenges and measure adoption (nemec2017measuring, ; kranch2015upgrading, ; holz2015tls, ). Collecting X.509 certificates through scans, 28 papers have shed light on the WebPKI prior to the widespread adoption of Certificate Transparency (vandersloot2016towards, ). There is also a large contingent of papers that have measured the Internet exposure of IoT devices (25 papers), ICS (14 papers), and security-relevant services (12 papers). Beyond understanding deployment patterns, a number of papers have been able to identify real-world attacks (durumeric2015neither, ), to reverse engineering attacker infrastructure (antonakakis2017understanding, ; marczak2015pay, ; marczak2014governments, ), and to conduct large-scale security notifications using ZMap data (li2016exploring, ; durumeric2014matter, ).
More networking-focused papers cover topics like DNS (24 papers), BGP/RPKI (12 papers), censorship (14 papers), and IP usage and NAT (10 papers). In addition to these studies, 53 other papers reference following our recommended practices when conducting other measurements. We provide a comprehensive breakdown of topics in Appendix B.
@@
Figure 4. ZMap by Country—The ten countries that emanate the most scan traffic by packet have varied ZMap usage.
@@ 2.3. Industry Adoption
Despite the plethora of publications, academic institutions are not responsible for most ZMap traffic. None of the top 100 ASes that emit the most ZMap traffic belong to universities; rather, most traffic originates from security companies and cloud providers. The network responsible for by far the most traffic is Google Cloud, which we find is predominantly used by Palo Alto Networks to power their Xpanse product.
To understand industry adoption, we categorized the organizations identified by Greynoise as using ZMap. We identify several categories of security products that are built on ZMap and fast Internet-wide scanning:
Attack Surface Management. With the shift to cloud-based infrastructure and a rise of ransomware attacks against enterprise services (e.g., MoveIt and VMWare ESXi), there has been an increased demand for companies to understand their Internet-facing infrastructure. Palo Alto Xpanse, Microsoft RiskIQ, and Rapid7 insightVM, along with more than a dozen other smaller companies, use ZMap as the basis for providing “attack surface management” products that give enterprises up-to-date data about their Internet-exposed risks and potentially unknown assets (“Shadow IT”).
Third-Party Risk Management. Building on the observation that externally visible security configuration and patching patterns correlate with data breaches and compromise (liu2015cloudy, ; zhang2014mismanagement, ; liu2015predicting, ), companies such as BitSight and FICO use ZMap to build security ratings that enable companies to understand the security of their supply-chain dependencies. SecurityScorecard, another major vendor, uses Masscan (ssc-masscan, ).
Internet Intelligence. More generically, several non-profit and commercial organizations use ZMap to collect data about IP addresses, threat actors, and Internet services, including BinaryEdge, Censys, IPInfo, and ShadowServer. Through these data sources, multiple countries now proactively monitor and notify organizations about risks (e.g., U.K. (ncsc-censys, )).
@@ 2.4. Malicious Use
Most security tools have the potential for both helping defenders and being misused by attackers. ZMap is no exception and there is evidence that ZMap has been used maliciously. While a significant proportion of scan traffic comes from networks with strong reputations (e.g., Google Cloud), attackers have also used “bulletproof” hosting providers to carry out scans for vulnerable services, including MSSQL, RDP, and Mikrotik’s router API (durumeric2014internet, ; anand2023aggressive, ). Attackers also have begun to repurpose ZMap to carry out DOS attacks (russia-dos, ), and, recently, two IoT botnets incorporated ZMap into their malware: in 2021, Palo Alto Networks found a Mirai variant adopted ZMap to search for SSH services (mirai-zmap, ), and, in 2023, a Medusa variant adopted ZMap to find Telnet services (medusa-zmap, ).
@@ 3. Lessons in Internet Scanning
L4 vs. L7 Discrepancies. Several studies have noted significant discrepancies between L4 and L7 responsiveness (mirian2016internet, ; durumeric2013analysis, ; holz2015tls, ; springall2016ftp, ; heninger2012mining, ; perino2018proxytorrent, ). Izhikevich et al. showed that TCP liveness does not reliably indicate the presence of an application-layer service because of pervasive middlebox deployment (izhikevich2021lzr, ). ZMap’s design also encouraged what Hiesgen et al. term “two-phase scanning” (hiesgen2022spoki, ), in which L4 service discovery and L7 service interrogation are performed separately. In response, some hosts “shun” two-phase scanners, exacerbating the perceived differences between L4 and L7 results (izhikevich2021lzr, ). Sattler et al. later devised a method for more accurately identifying highly L4-but-not-L7 responsive prefixes (sattler2023packed, ). These differences fundamentally limit ZMap’s utility (as a standalone L4 tool) to discovering potential services, requiring most work to be completed in follow-up L7 scans and shifting our focus to downstream tools like LZR (izhikevich2021lzr, ) and ZGrab (zgrab2, ).
Visibility and Consistency. One challenge of running Internet scans is lack of ground truth for validation. Wan et al. showed that the ZMap paper slightly overestimated the visibility achieved by the tool, and that a single-probe scan actually misses about 2.7% of HTTP(S) hosts (wan2020origin, ). As can be seen in Figures 1 of Hastings et al. (hastings2016weak, ) and Chung et al. (chung2016measuring, ), different organizations using ZMap sometimes see different results. However, in even the most egregious cases (e.g., Censys (durumeric2015search, )), vantage points only miss under 5% of services due to blocking (in addition to transient drop); the bulk of loss is typically driven by a handful of small service and cloud providers (wan2020origin, ).
We now recommend that the best way to mitigate transient drop is to scan from 2–3 geographically and topologically diverse vantages, rather than to send multiple probes from a single scanner, since both probes are oftentimes lost. Results can also differ across scanning tools: Adrian et al. showed that, despite following a similar high-level approach, Masscan (graham2014masscan, ) finds notably fewer hosts than ZMap, likely due to biases in Masscan’s randomization algorithm (adrian2014zippier, ).
@@ 4. ZMap Codebase
When we released ZMap, we had little idea what community, if any, would emerge. Over the past ten years, more than 80 individuals have committed code to ZMap—we are deeply thankful to those contributors. Despite the high number of committers, 90% of ZMap code has been written by five individuals. Most contributions to the code have been made by industry researchers and software developers. Academic involvement has been limited: of the 11 external contributors who changed more than 100 LoC, only one is an academic researcher. Academic groups most frequently contribute probe modules or bugfixes; when academic groups made improvements to core functionality, improvements tend to be forked and renamed (e.g., XMap (xmap, ) and ZMapv6 (gasser2016scanning, ), which implemented the same IPv6 functionality) rather than upstreamed.
Beyond usability improvements and bug fixes, there have been several significant changes in how ZMap operates that make the original paper’s description of the tool out of date. We describe the most significant of these changes below.
@@ 4.1. Address and Port Generation
Hosts vs. Targets. While ZMap originally tracked statistics by IP address, our randomization approach selects from a pool of (IP, port) “targets”, rather than considering IPs and ports independently. As a result, options like “max hosts” cannot be supported. All results, statistics, and configuration in ZMap are now based on the notion of an IP/port target.
@@
Identifying Generators. To create a new permutation of the address space for every scan, ZMap originally identified a random generator (i.e., primitive root) of the appropriate multiplicative group by identifying a generator of ($\mathbb{Z}_{p-1}$, +) and then mapping it into ($\mathbb{Z}^{*}_{p}$, $\times$)  (durumeric2013zmap, ). This was practical because a generator of the additive group is any integer coprime with $p-1$. This can be efficiently tested against randomly drawn integers, and there are $\phi(2^{32}+14)\approx 10^{9}$ generators of the additive group, resulting in an average of four attempts to identify a generator. Nearly all additive generators could be mapped into usable generators in the multiplicative group, since the only constraint on the multiplicative generators was that they were less than $2^{32}$ (to ensure safe multiplication using 64-bit arithmetic).
@@
For multiport scans, to support iterating over $2^{48}$ elements, we need to efficiently find multiplicative generators smaller than $2^{16}$ in a $2^{48}$ search space. Unfortunately, because an additive generator can map to a multiplicative generator anywhere in the group, only $1/2^{32}$ candidate additive generators are usable. To address this, we flipped our approach. For each group defined by prime $p$, we precalculate and store the prime factorization of $p-1=k_{1}^{a_{1}}k_{2}^{a_{2}}\dots k_{n}^{a_{n}}$ for distinct primes $k_{i}\,(1\leq i\leq n)$. At runtime, we generate a random candidate generator $g\in[2\mathrel{{.}\,{.}}\nobreak 2^{16}-1]$ that is guaranteed to keep future arithmetic within the 64-bit address space. Then we ensure that $g$ is a generator of $\mathbb{Z}^{*}_{p}$ by checking that: $g^{(p-1)/k_{i}}\mod p\neq 1\,\forall\,i\in[1\mathrel{{.}\,{.}}\nobreak n]$. Similarly to the first approach, this requires an average of four attempts.
@@
Figure 5. Sliding Window Duplicate Rate—We moved to a sliding window-based approach for deduplicating responses as part of supporting multiple ports. A window size of at least $10^{7}$ eliminates nearly all duplicates.
@@
Response Deduplication. Hosts frequently send back repeated responses, in some cases indefinitely. While we initially thought these were due to broken TCP implementations, Goldblatt et al. showed that some hosts will aggressively send tens of thousands of response packets in response to a SYN packet, which they term “blowback” (goldblatt2023blowback, ). We originally filtered out duplicate responses using a paged $2^{32}$-bit bitmap, which used 512 MB of memory. While this approach guarantees no duplicates, extending it to the 48-bit space of IPs and ports would require 35 TB. Instead, we switch to maintaining a sliding window of the last $n$ IP/Port responses, using a Judy array (baskins2000judy, ). As can be seen in Figure 5, maintaining a window of $10^{7}$ entries is effective to filter nearly all duplicate responses, and lower scan rates can make do with smaller window sizes. We found zero duplicates for the three trials of 1 Gbps scans targeting TCP/80 in April 2024.
4.2. Sharding

### v2

@@ Abstract
Since ZMap’s debut in 2013, networking and security researchers have used the open-source scanner to write hundreds of research papers that study Internet behavior. In addition, ZMap has been adopted by the security industry to build new classes of enterprise security and compliance products. Over the past decade, much of ZMap’s behavior—ranging from its pseudorandom IP generation to its packet construction—has evolved as we have learned more about how to scan the Internet. In this work, we quantify ZMap’s adoption over the ten years since its release, describe its modern behavior (and the measurements that motivated changes), and offer lessons from releasing and maintaining ZMap for future tools.
@@ 1. Introduction
In 2013, Durumeric et al. released ZMap (durumeric2013zmap, ), an open-source network scanner that made it dramatically easier to scan the entire IPv4 address space. Since then, more than 300 research papers have used ZMap to uncover protocol flaws (adrian2015imperfect, ; aviram2016drown, ), shed light on the WebPKI (durumeric2013analysis, ), reverse engineer mercenary spyware (marczak2014governments, ), understand headline events like Heartbleed (durumeric2014matter, ) and Mirai (antonakakis2017understanding, ), and more. Beyond research, security companies have developed products on top of ZMap to continuously monitor organizations’ attack surfaces and their third-party dependencies. At the same time, like many security tools, ZMap has also been adopted by attackers to identify vulnerable systems. In aggregate, ZMap now accounts for over one-third of all Internet-wide scan traffic.
Since its initial release, ZMap has evolved as we learned more about how to scan the Internet and better understood researchers’ needs. Yet, with the exception of Adrian et al.’s write up of ZMap’s 10GbE re-architecture (adrian2014zippier, ), the project team has not documented these changes in the research literature. Of the improvements Adrian et al. introduced, both the lock-free randomization algorithm and fast packet transmission approach have since changed. As we maintained and improved ZMap, we also learned a great deal about how to (and not to) build Internet measurement tools.
Motivated by requests from the community to document what we have learned (claffy2019workshop, ), in this work, we present a retrospective analysis of ZMap. We cover how ZMap has been adopted (§2), advances in Internet scanning that challenged the tool’s initial design assumptions (§3), significant changes to ZMap’s behavior (§4), and what we learned maintaining and improving ZMap (§5). We hope that by analyzing this endeavor, we can help the researchers who build the next generation of Internet measurement tools.
@@ 2. Usage and Internet Landscape
Since ZMap’s release in 2013, its adoption has steadily increased (durumeric2014internet, ; anand2023aggressive, ). Today, over 33% of all Internet-wide IPv4 scan traffic can be fingerprinted as coming from ZMap. In this section, we provide a high-level overview of ZMap’s adoption by academic researchers, security companies, and malicious actors.
Figure 1. ZMap-Attributed TCP Scan Traffic—ZMap growth has accelerated significantly since 2020. In Q1 2024, 35% of Internet-wide IPv4 TCP scan traffic (by packet) came from ZMap.
@@ 2.1. Empirical Analysis
One year after ZMap’s release, in 2014, Durumeric et al. found relatively little adoption and that ZMap was primarily used to study academically interesting protocols like HTTP(S) (durumeric2014internet, ). More recently, in 2021, Anand et al. noted that many “aggressive” scans were from ZMap or Masscan (anand2023aggressive, ). Using the same methodology as these two studies (durumeric2014internet, ; anand2023aggressive, ), we measure ZMap adoption by analyzing scans that target at least ten IPs in the ORION Network Telescope (orion, ). Our analysis is limited to TCP scans, since ORION identifies TCP, ICMP and UDP scanning flows, but only tags scanning tools for TCP flows. In addition, we note that forks of ZMap that remove the static identifying IP ID of 54321 will not be attributed to ZMap.
ZMap usage has increased dramatically over the past four years (Figure 1). From January 1 to March 31, 2024, 35.4% of all Internet-wide IPv4 TCP scan packets originated from ZMap. While it is possible that ICMP and UDP scans use other tools, the vast majority of all scan traffic is TCP-based and ZMap still accounts for 33% of all scan packets with only TCP traffic attributed. ZMap scans follow a different distribution of targeted ports than global scans (Figures 2 and 3). For example, ZMap accounts for only 12% of TCP/23 but 69% of TCP/80 and 73% of TCP/8080. In the most extreme case, 99.5% of traffic targeting TCP/8728 (MikroTik router API) globally is from ZMap, driving it to the sixth most scanned port globally. There are also dramatic regional differences in ZMap adoption (Figure 4). For example, while ZMap accounts for more than 66% of scan packets from U.S. hosts, less than 0.5% of Russian scan packets are from ZMap. The outsized ZMap use in the U.S. is largely due to adoption by American security companies (§2.3). In parallel to our work, Griffioen et al. conducted a more in-depth analysis of how scanning has evolved over the past decade (scanning-ten, ); the study found that 59% of Internet-wide scans in 2024 used ZMap.
Figure 2. All TCP Scans (Top Ports by Packet)
Figure 3. ZMap Scans (Top Ports By Packet)
@@ 2.2. Academic Research
ZMap has been used for a vast range of research purposes, from showing the possible compromise of RSA keys through transient faults (sullivan2022open, ) to measuring NAT64 deployment (hsu2024first, ). To understand the studies that ZMap has enabled, one author manually analyzed 1,034 papers that cite or reference ZMap through April 2024 via Google Scholar and categorized papers using thematic analysis (braun2012thematic, ). We exclude dissertations (since these are often comprised of published papers) as well as studies that used Censys (durumeric2015search, ). In total, we identified 307 research papers directly based on ZMap data.
While ZMap is a general measurement tool, it has most prominently been used by the security community. Notably, ZMap has been used in 38 papers to uncover protocol weaknesses in TLS and underlying cryptographic primitives (beurdouche2015messy, ; adrian2015imperfect, ; aviram2016drown, ) as well as to uncover deployment challenges and measure adoption (nemec2017measuring, ; kranch2015upgrading, ; holz2015tls, ). Collecting X.509 certificates through scans, 28 papers have shed light on the WebPKI prior to the widespread adoption of Certificate Transparency (vandersloot2016towards, ). There is also a large contingent of papers that have measured the Internet exposure of IoT devices (25 papers), ICS (14 papers), and security-relevant services (12 papers). Beyond understanding deployment patterns, a number of papers have been able to identify real-world attacks (durumeric2015neither, ), to reverse engineering attacker infrastructure (antonakakis2017understanding, ; marczak2015pay, ; marczak2014governments, ), and to conduct large-scale security notifications using ZMap data (li2016exploring, ; durumeric2014matter, ). We encourage the security community to embrace research that builds Internet measurement tools and techniques since these are frequently used to understand and improve security practices.
More networking-focused papers cover topics like DNS (24 papers), BGP/RPKI (12 papers), censorship (14 papers), and IP usage and NAT (10 papers). In addition to these studies, 53 other papers reference following the recommended practices when conducting other measurements. We provide more details in Appendix B.
@@
Figure 4. ZMap by Country—The ten countries that emanate the most Internet scan traffic by packet have varied ZMap usage.
Despite the plethora of publications, academic networks are not responsible for most ZMap traffic, likely because research experiments do not require continuous long-term scanning of a large number of ports. None of the top 100 ASes that emit the most ZMap traffic belong to universities; rather, most traffic originates from security companies and cloud providers. For example, the provider responsible for—by far—the most ZMap scan traffic is Google Cloud (GCP). Examining the reverse DNS records for scanning IPs, we find that GCP is predominately used by Palo Alto Networks to power their Xpanse Attack Surface Management product.
@@ 2.3. Industry Adoption
To understand industry adoption, we categorized the organizations identified by Greynoise as using ZMap and mapped these to broad industry categories of security products (e.g., as defined by Gartner):
Attack Surface Management. With the shift to cloud-based infrastructure and a rise of ransomware attacks against enterprise services (e.g., MoveIt and VMWare ESXi), there has been an increased demand for companies to understand their Internet-facing infrastructure. Palo Alto Xpanse, Microsoft RiskIQ, and Rapid7 insightVM, along with numerous other smaller companies, use ZMap as the basis for providing “attack surface management” products that give enterprises up-to-date data about their Internet-exposed risks and potentially unknown assets.
Third-Party Risk Management. Building on the observation that externally visible security configuration and patching patterns correlate with data breaches and compromise (liu2015cloudy, ; zhang2014mismanagement, ; liu2015predicting, ), companies such as BitSight and FICO use ZMap to build security ratings that enable companies to understand their supply-chain security.
Internet Intelligence. Several non-profits and companies use ZMap to collect data about IP addresses, threat actors, and Internet services, including BinaryEdge, Censys, IPInfo, and ShadowServer. Using this data, multiple countries proactively monitor for and notify organizations about risks (e.g., U.K. (ncsc-censys, )).
@@ 2.4. Malicious Use
Most security tools have the potential for both helping defenders and being misused by attackers. ZMap is no exception and there is evidence that ZMap has been used maliciously. While it is difficult to ascertain the intent of network traffic from a shared cloud provider without any application-layer visibility (hiesgen2022spoki, ; izhikevich2023cloud, ), past darknet analysis has shown that attackers have used “bulletproof” hosting providers to carry out scans for vulnerable services, including MSSQL, RDP, and Mikrotik’s router API (durumeric2014internet, ; anand2023aggressive, ). Anecdotally, attackers also have begun to repurpose ZMap to carry out DOS attacks (russia-dos, ), and, recently, two IoT botnets incorporated ZMap into their malware: in 2021, Palo Alto Networks found a Mirai variant adopted ZMap to search for SSH services (mirai-zmap, ), and, in 2023, a Medusa variant adopted ZMap to find Telnet services (medusa-zmap, ).
@@ 3. Lessons in Internet Scanning
L4 vs. L7 Discrepancies. Several studies have noted significant discrepancies between L4 and L7 responsiveness (mirian2016internet, ; durumeric2013analysis, ; holz2015tls, ; springall2016ftp, ; heninger2012mining, ; perino2018proxytorrent, ). Izhikevich et al. showed that TCP liveness does not reliably indicate service presence because of pervasive middlebox deployment (izhikevich2021lzr, ). ZMap’s design also encouraged what Hiesgen et al. term “two-phase scanning” (hiesgen2022spoki, ), in which L4 service discovery and L7 service interrogation are performed separately. In response, some hosts “shun” two-phase scanners, exacerbating the perceived differences between L4 and L7 results (izhikevich2021lzr, ). Sattler et al. later devised a method for more accurately identifying highly L4-but-not-L7 responsive prefixes (sattler2023packed, ). These differences fundamentally limit ZMap’s utility (as a standalone L4 tool) to discovering potential services, requiring most work to be completed in follow-up L7 scans and shifting our focus to downstream tools like LZR (izhikevich2021lzr, ) and ZGrab (zgrab2, ).
Visibility and Consistency. One challenge of running Internet scans is the lack of ground truth for validation. Wan et al. showed that the ZMap paper slightly overestimated the visibility achieved by the tool, and that a single-probe scan actually misses about 2.7% of HTTP(S) hosts (wan2020origin, ). Figures 1 of Hastings et al. (hastings2016weak, ) and Chung et al. (chung2016measuring, ) show that different organizations using ZMap sometimes see different results. However, in even the most egregious cases (e.g., Censys (durumeric2015search, )), vantage points miss under 5% of services due to blocking; the bulk of loss is typically driven by a handful of small service and cloud providers (wan2020origin, ). For those who need comprehensive coverage, Wan et al. recommends that the best way to mitigate transient drop is to scan from 2–3 geographically and topologically diverse vantages, rather than to send multiple probes from a single scanner, since both probes are oftentimes lost. Results can also differ across scanning tools: Adrian et al. showed that, despite following a similar high-level approach, Masscan (graham2014masscan, ) finds notably fewer hosts than ZMap, likely due to biases in its randomization algorithm (adrian2014zippier, ).
@@ 4. ZMap Codebase
When we released ZMap, we had little idea what community, if any, would emerge. Over the past ten years, more than 80 individuals have committed code to ZMap—we are deeply thankful to those contributors. Despite the high number of committers, 90% of ZMap code has been written by five individuals. Most contributions to the code have been made by industry and academic involvement has been limited: of the 11 external contributors who changed more than 100 LoC, only one is an academic researcher. Academic groups most frequently contributed probe modules or bugfixes; when academics made improvements to core functionality, improvements tend to be forked and renamed (e.g., XMap (xmap, ) and ZMapv6 (gasser2016scanning, ), which implemented the same IPv6 functionality) rather than upstreamed. Funding from the NSF Internet Measurement Research: Methodologies, Tools, and Infrastructure (IMR) program (nsf-imr, ) has been critical in supporting ZMap’s continued development.
Beyond usability improvements and bug fixes, there have been several fundamental changes in how ZMap operates beyond the original paper’s description. We describe these changes below.
@@ 4.1. Address and Port Generation
Hosts vs. Targets. ZMap originally tracked statistics by IP address. The new randomization approach selects from a pool of (IP, port) “targets”, rather than considering IPs and ports independently. As a result, all data, metadata, and configuration in ZMap are now based on the notion of an (IP, port) target. This enables randomization across the IP–Port space (e.g., rather than rotating scans across ports with each port operating independently), but precludes options like “max hosts” without significant additional state.
@@
Identifying Generators. To create a new permutation of the address space for every scan, ZMap originally identified a random generator (i.e., primitive root) of the appropriate multiplicative group by identifying a generator of ($\mathbb{Z}_{p-1}$, +) and then mapping it into ($\mathbb{Z}^{*}_{p}$, $\times$)  (durumeric2013zmap, ). This was practical because a generator of the additive group is any integer coprime with $p-1$, which is efficiently testable with the Euclidean algorithm against randomly drawn integers. There are $\phi(2^{32}+14)\approx 10^{9}$ generators of the additive group, resulting in an average four attempts to identify a generator. Nearly all additive generators could be mapped into usable generators in the multiplicative group, since the only constraint on the multiplicative generators was that they were less than $2^{32}$ (to ensure safe multiplication using 64-bit arithmetic).
@@
For multiport scans, to support iterating over $2^{48}$ elements, we need to efficiently find multiplicative generators smaller than $2^{16}$ in a $2^{48}$ search space. Unfortunately, because an additive generator can map to a multiplicative generator anywhere in the group, only $1/2^{32}$ candidate additive generators are usable. To address this, we flipped our approach. For each group defined by prime $p$, we precalculate and store the prime factorization of $p-1=k_{1}^{a_{1}}k_{2}^{a_{2}}\dots k_{n}^{a_{n}}$ for distinct primes $k_{i}\,(1\leq i\leq n)$. At runtime, we generate a random candidate generator $g\in[2\mathrel{{.}\,{.}}\nobreak 2^{16}-1]$ that is guaranteed to keep future arithmetic within the 64-bit address space. Then we ensure that $g$ is a generator of $\mathbb{Z}^{*}_{p}$ by checking that: $g^{(p-1)/k_{i}}\mod p\neq 1\,\forall\,i\in[1\mathrel{{.}\,{.}}\nobreak n]$. This requires an average four attempts.
@@
Figure 5. Sliding Window Duplicate Rate—We moved to a sliding window approach for deduplicating responses to support multiple ports. A window size of $10^{6}$ eliminates nearly all duplicates.
@@
Response Deduplication. Hosts frequently send back repeated responses, in some cases indefinitely. While we initially thought these were due to broken TCP implementations, Goldblatt et al. showed that some hosts will aggressively send tens of thousands of response packets, which they term “blowback” (goldblatt2023blowback, ). We originally filtered out duplicate responses using a paged $2^{32}$-bit bitmap, which used 512 MB of memory. While this approach guarantees no duplicates, extending it to the 48-bit space of IPs and ports would require 35 TB. Instead, we switch to maintaining a sliding window of the last $n$ IP/Port responses, using a Judy array (baskins2000judy, ). As can be seen in Figure 5, a window of $10^{6}$ entries (ZMap default) is effective to filter nearly all duplicate responses, and lower scan rates can make do with smaller window sizes. We found zero duplicates for three trials of 1 Gbps scans targeting TCP/80 in April 2024.
4.2. Scan Sharding